OpenStudy (anonymous):

use matrices A and B. Compute B-A, if you can. A= -5 4 B= -2 7 -3 -8 2 1 -6 0

OpenStudy (anonymous):

@whpalmer4

OpenStudy (phi):

they have to be the same shape to add or subtract matrices.

OpenStudy (anonymous):

so is it not possible?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[\left( \begin{array}{ccc} -2 & 7 & -3 \\ 1 & -6 & 0 \\ \end{array} \right)-\left( \begin{array}{cc} -5 & 4 \\ -8 & 2 \\ \end{array} \right)\] How would you propose to do that? Remember, \[\left( \begin{array}{ccc} a & b \\ c & d \\ \end{array} \right)-\left( \begin{array}{cc} e & f \\ g & h \\ \end{array} \right)\] is just another way of writing\[a-e\]\[b-f\]\[c-g\]\[d-h\]What would we do with the extra (or missing, depending on how you look at it) numbers?

OpenStudy (anonymous):

so not possible correct?

OpenStudy (whpalmer4):

Yes, not possible, unless you invent some mathematics to make it possible.